OpenStudy (anonymous):

Find the derivative of the following from first principle: a) f(x) = x-x^3 B) y = (2x-5)^2

OpenStudy (anonymous):

if your looking for the derivative (y') a. f(x) = x - x^3 = 1 - 3x^2 b. y = (2x - 5)^2 [using chain rule] 2(2x - 5) (2) =4(2x - 5) [not sure about "a" though ^^;;]

OpenStudy (anonymous):

I tried using the rule but I keep getting the wrong answer through the method

OpenStudy (anonymous):

This rule I meant : f(x+h)-f(x)/h

OpenStudy (shubhamsrg):

whats your attempt ?

OpenStudy (anonymous):

k got it.

OpenStudy (anonymous):